What total capacitances can you make by connecting a 4.12 µF and 7.72 µF capacitor together?

series     µF
parallel     µF
Answer #1

When these capacitor are connected in series then

\frac{1}{C_s}= \frac{1}{C_1}+\frac{1}{C_2}

\frac{1}{C_s}= \frac{1}{4.12 \mu F}+\frac{1}{7.72 \mu F}

C_s= 2.69 \mu F

When connected in parallel

C_P= C_1+ C_2 = 4.12 \mu F + 7.72 \mu F\Rightarrow 11.84 \mu F